# Machine-learned flow estimation with sparse data — exemplified for the rooftop of a UAV vertiport

### v1

@@ IV.2 Flow estimation on the rooftop of a UAV vertiport from sensor signals
To gain a comprehensive understanding of the framework, we compare $5$ cases with different model structures and data categories. Table 4 displays the mean estimation error for the various combinations within the framework. The results are obtained for the velocity sensor; for the pressure sensor, see Appendix B. Here, the full manifold refers to using all velocity field data $\bm{u}^{*m}$ in the modeling process (standard ISOMAP), while the coarse-grained manifold indicates using the cluster centroids $\bm{c}^{*}_{k}$ (cISOMAP). The training data utilizes the existing database, with single $U_{\infty}$ set at $b_{4}=7.9$ m/s and multiple $U_{\infty}$ selected randomly. The testing data is gathered from additional simulations under random wind conditions, with interpolation covering wind speeds from $b_{0}$ (0.2 m/s) to $b_{8}$ (20.7 m/s) and extrapolation from $U_{\infty}=20.7$ m/s to $U_{\infty}=30$ m/s. For the training data, the nearest neighbor for the wind condition estimator is $K_{w}=1$, while for the testing data, $K_{w}=2$. The comparison between the full manifold and the coarse-grained manifold evaluates the impact of the clustering approach, while comparisons across different data sources reveal training and generalization errors. Additionally, comparing non-dimensionalized multiple wind speeds with a single wind speed validates Reynolds number independence, and extrapolation tests performance under rare wind conditions. Comparing Case $1$ and Case $2$, model accuracy with multiple non-dimensionalized wind speeds is the same as that with a single speed (all 00 % error), indicating that our framework corroborates Reynolds number independence. This independence simplifies the manifold structure and ensures consistent performance across varying wind speeds. The comparison of Case $2$ and Case $3$ shows a $0.55\%$ increase in error with coarse-graining. Despite this increase, clustering typically reduces computational load by one order of magnitude based on the number of clusters. This substantial computational efficiency makes cISOMAP particularly valuable for handling large-scale datasets, emphasizing its practical benefits in industrial applications, where computational power and time are critical constraints. Comparing Case $3$ and Case $4$ shows that the generalization error remains within acceptable limits. The observed error primarily arises from biases introduced by the wind condition estimator rather than from the manifold model itself. The increase in error also indicates that improvements in the estimator could further enhance overall accuracy. Finally, the comparison between Case $4$ and Case $5$ demonstrates that the framework is highly resilient to extrapolation, which also validates Reynolds number independence. The minimal additional estimation error observed when extrapolating to rare wind conditions highlights the robustness of the model, ensuring reliable predictions even under less common scenarios. This resilience reveals the practical value of the framework in real-world applications where encountering extreme conditions is essential for further operations.
Figure 14: Estimation error of Case $5$: The estimation error of the oncoming velocity $\hat{U}_{\infty}$ is shown in (a) for wind speed $U_{\infty}$ and in (b) for wind angle $\alpha$. The normalized squared error (representation error) of the velocity field is presented in (c) against wind speed $U_{\infty}$ and in (d) against wind angle $\alpha$. The normalized mean squared error is highlighted in red ($E=6.55\%$), while the blue dot indicates the snapshot for further visualization.
Figure 15: Estimated velocity field using the velocity sensor: The top right shows the slice locations, and the dashed line in the flow visualization indicates the position of the vertiport. Top: CFD dataset. Bottom: Estimated field. Snapshot wind conditions: $U_{\infty}=26.55$ m/s, $\alpha=65.22^{\circ}$, with an estimation error of $E=2.25\%$.
Figure 14 shows the estimation error of $\hat{U}_{\infty}$ and $\hat{u}$ for all test wind conditions in Case $5$. The error distribution is relatively uniform across various wind speeds and angles, indicating no specific conditions with significantly higher errors, which demonstrates the generalization capability of the framework. The most significant error among all estimated velocity fields is within $10\%$, while most have errors below $5\%$. Even for the extrapolated wind conditions which are not included in the CFD database, the error distribution remains consistent with the rest of the dataset. An estimated snapshot, marked in blue and close to the normalized mean squared error, is selected for flow visualization in Figure 15. This visualization includes the estimated flow field across four rooftop transverse planes at heights of $z=0.2$ m, $0.4$ m, $0.6$ m, and $0.8$ m. The accurate prediction of the wind angle is evident, with the flow structure in the estimated data closely resembling that of the CFD data, highlighting the framework’s effectiveness in capturing full-state flow dynamics.
@@ V Conclusions
The proposed framework approximates the mean flow data from many operating conditions as a non-dimensionalized coarse-grained manifold. The first encoder normalizes the flow of data with the oncoming wind speed. The second encoder projects this non-dimensionalized data into a least-order manifold via ISOMAP. This manifold is coarse-grained with cluster-based collocation points to reduce the computational load of the en- and decoding. The decoders reverse this process: $k$NN performs the initial decoding, followed by a second decoder that restores the normalization. This manifold simplifies the flow estimation from sensor signals. First, these signals are employed to estimate the oncoming wind speed. We emphasize that the assumed Reynolds-number independence allows us to infer wind speeds far beyond the database. Second, the correspondingly non-dimensionalized sensor signals infer the latent variables on the manifold. These latent variables are decoded into full-state flow fields as described above. Reynolds number independence enables this methodology to extrapolate wind conditions outside the training set, a feature that traditional data-centric frameworks cannot achieve.
We demonstrate the effectiveness of this framework for the flow on the rooftop of a vertiport at Reynolds number from $4\times 10^{4}$ to $6\times 10^{6}$. Flows with 1080 different oncoming velocities and angles are simulated with Reynolds averaged Navier-Stokes simulations for the same inlet boundary-layer layer. Hence, the flow is expected to lie on a two-dimensional manifold with two latent variables representing the speed and angle of the oncoming flow. Intriguingly, the non-dimensionalized manifold learner corroborates Reynolds number independence: A limit-cycle manifold leads to a representation error of about $2\%$, i.e. the non-dimensionalized flow is effectively a function of wind direction.
@@
Summarizing, the proposed estimation framework provides a robust, interpretable, and generalizable approach to flow estimation and modeling. The results not only enhance our understanding of urban wind fields but also open avenues for applications such as gust-safe drone delivery and flight control. Future work aims to extend this estimation framework to predict the wind for drone trajectories in Shenzhen. These efforts including multiple building clusters, variable wind profiles and flow instabilities.
@@
Figure 18: Same as figue 10, but for the building cluster.
@@ Appendix B Flow estimation by the pressure sensor
Case Error $\hat{U}_{\infty}$ (%) Error (%)
@@
Figure 21: Same as figure 15, but from the pressure sensor, with $E=2.35\%$.

### v2

@@ IV.3 Error sources of the model
To gain a comprehensive understanding of the framework, we compare $5$ cases with different model structures and data categories. Table 4 displays the mean estimation error for the various combinations within the framework. The results are obtained for the velocity sensor; for the pressure sensor, see Appendix B. Here, the full manifold refers to using all velocity field data $\bm{u}^{*m}$ in the modeling process (standard ISOMAP), while the coarse-grained manifold indicates using the cluster centroids $\bm{c}^{*}_{k}$ (cISOMAP). The training data utilizes the existing database, with single $U_{\infty}$ set at $b_{4}=7.9$ m/s and multiple $U_{\infty}$ selected randomly. The testing data is gathered from additional simulations under random wind conditions, with interpolation covering wind speeds from $b_{0}$ (0.2 m/s) to $b_{8}$ (20.7 m/s) and extrapolation from $U_{\infty}=20.7$ m/s to $U_{\infty}=30$ m/s. For the training data, the nearest neighbor for the wind condition estimator is $K_{w}=1$, while for the testing data, $K_{w}=2$. The comparison between the full manifold and the coarse-grained manifold evaluates the impact of the clustering approach, while comparisons across different data sources reveal training and generalization errors. Additionally, comparing non-dimensionalized multiple wind speeds with a single wind speed validates Reynolds number independence, and extrapolation tests performance under rare wind conditions. Comparing Case $1$ and Case $2$, model accuracy with multiple non-dimensionalized wind speeds is the same as that with a single speed (all 00 % error), indicating a low training error from the full manifold. This independence simplifies the manifold structure and ensures consistent performance across varying wind speeds. The comparison of Case $2$ and Case $3$ shows a $0.55\%$ increase in error with coarse-graining. Despite this increase, clustering typically reduces computational load by one order of magnitude based on the number of clusters. This substantial computational efficiency makes cISOMAP particularly valuable for handling large-scale datasets, emphasizing its practical benefits in industrial applications, where computational power and time are critical constraints. Comparing Case $3$ and Case $4$ shows that the generalization error remains within acceptable limits. The observed error primarily arises from biases introduced by the wind condition estimator rather than from the manifold model itself. The increase in error also indicates that improvements in the estimator could further enhance overall accuracy. Finally, the comparison between Case $4$ and Case $5$ demonstrates that the framework is highly resilient to extrapolation, which also validates Reynolds number independence. The minimal additional estimation error observed when extrapolating to rare wind conditions highlights the robustness of the model, ensuring reliable predictions even under less common scenarios. This resilience reveals the practical value of the framework in real-world applications where encountering extreme conditions is essential for further operations.
@@ V Conclusions
The proposed framework approximates the mean flow data from many operating conditions as a non-dimensionalized coarse-grained manifold. The first encoder normalizes the flow data with the oncoming wind speed. The second encoder projects this non-dimensionalized data into a least-order manifold via ISOMAP. This manifold is coarse-grained with cluster-based collocation points to reduce the computational load of the en- and decoding. The decoders reverse this process: $k$NN performs the initial decoding, followed by a second decoder that restores the normalization. This manifold simplifies the flow estimation from sensor signals. First, these signals are employed to estimate the oncoming wind speed. We emphasize that the assumed Reynolds-number independence allows us to infer wind speeds far beyond the database. Second, the correspondingly non-dimensionalized sensor signals infer the latent variables on the manifold. These latent variables are decoded into full-state flow fields as described above. Reynolds number independence enables this methodology to extrapolate wind conditions outside the training set, a feature that traditional data-centric frameworks cannot achieve.
We demonstrate the effectiveness of this framework for the flow on the rooftop of a vertiport at Reynolds number from $4\times 10^{4}$ to $6\times 10^{6}$. Flows with 1080 different oncoming velocities and angles are simulated with Reynolds averaged Navier-Stokes simulations for the same inlet boundary layer. Hence, the flow is expected to lie on a two-dimensional manifold with two latent variables that exhibit a strong correlation with the angle of the oncoming flow. Intriguingly, the non-dimensionalized manifold learner corroborates Reynolds number independence: a limit-cycle manifold leads to a representation error of about $2\%$, i.e., the non-dimensionalized flow is effectively a function of wind direction.
@@
Summarizing, the proposed estimation framework provides a robust, interpretable, and generalizable approach to flow estimation and modeling. The results not only enhance our understanding of urban wind fields but also open avenues for applications such as gust-safe drone delivery and flight control. Future work aims to extend this estimation framework to predict the wind for drone trajectories in Shenzhen. These efforts include multiple building clusters, variable wind profiles and flow instabilities.
@@
Figure 18: Same as figure 10, but for the building cluster.
@@ Appendix B Flow estimation by the pressure sensor
Case $\hat{U}_{\infty}$ Error (%) Error (%)
@@
Figure 21: The same description of figure 15 applies, but from the pressure sensor, with $E=2.35\%$.
@@
Appendix C Robustness analysis
In industrial applications, input signals frequently contain noise. To assess the robustness of our framework, we superimpose increasing levels of white Gaussian noise, $\bm{n}(t)$, to the clean sensor signal and evaluate its performance. The noise level, ranging from $0\%$ to $40\%$, is defined as a proportion of instantaneous signal power:
$\mathrm{Noise\;level}=P_{\bm{n}}/P_{\bm{s}},$ (37)
where $P_{\bm{n}}$ is the noise power, and $P_{\bm{s}}$ is the clear signal power, for the $t$-th input signal:
$P_{\bm{n}}(t)=\bm{n}(t)^{\top}\bm{n}(t),$ (38)
$P_{\bm{s}}(t)=\bm{s}(t)^{\top}\bm{s}(t),$ (39)
The fourth testing dataset with the velocity sensor is used, and the results are shown in figure 22. As expected, there is a strong correlation between noise level and estimation error, indicating that the framework is sensitive to noise. Interestingly, the relationship is almost linear. Higher noise levels gradually reduce the model’s accuracy compared to the reference clean data. Nonetheless, the model retains acceptable accuracy when noise levels remain below $20\%$.
Figure 22: Estimation error with noisy input for the velocity sensor in testing dataset 4. The noisy sensor signal is created by adding white Gaussian noise of varying levels to the clean signal. As a reference, the estimation error of the clean (noise-free) signal is $1.972\%$.